# FedAIoT: A Federated Learning Benchmark for Artificial Intelligence of Things

### v2

@@ Abstract
There is a significant relevance of federated learning (FL) in the realm of Artificial Intelligence of Things (AIoT). However, most existing FL works do not use datasets collected from authentic IoT devices and thus do not capture unique modalities and inherent challenges of IoT data. To fill this critical gap, in this work, we introduce FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of IoT devices. These datasets cover unique IoT modalities and target representative applications of AIoT. FedAIoT also includes a unified end-to-end FL framework for AIoT that simplifies benchmarking the performance of the datasets. Our benchmark results shed light on the opportunities and challenges of FL for AIoT. We hope FedAIoT could serve as an invaluable resource to foster advancements in the important field of FL for AIoT. The repository of FedAIoT is maintained at https://github.com/AIoT-MLSys-Lab/FedAIoT.
@@ 1 Introduction
The proliferation of the Internet of Things (IoT) such as smartphones, drones, and sensors deployed at homes, and the gigantic amount of data they capture, have revolutionized the way we work, live, and interact with the world. The advances in Artificial Intelligence (AI) have boosted the integration of IoT and AI that turns Artificial Intelligence of Things (AIoT) into reality. However, data captured by IoT devices usually contain privacy-sensitive data. In recent years, federated learning (FL) has emerged as a privacy-preserving solution that enables extracting knowledge from the collected data while keeping the data on the devices (Kairouz et al., 2021; Wang et al., 2021).
Despite the significant relevance of FL, most existing FL works are conducted on well-known datasets. We present an overview of the datasets different FL papers have used from the pioneering works to more recent publications in Table 9 These datasets, however, do not originate from authentic IoT devices and thus fail to capture the unique modalities and inherent challenges associated with real-world IoT data. This discrepancy underscores a strong need for an IoT-oriented FL benchmark to fill this critical gap.
In this work, we present FedAIoT, an FL benchmark for AIoT (Figure 1). At its core, FedAIoT includes eight well-chosen datasets collected from a wide range of IoT devices from smartwatch, smartphone and Wi-Fi routers, to drones, smart home sensors, and head-mounted device that either have already become an indispensable part of people’s daily lives or are driving emerging applications. These datasets encapsulate a variety of unique IoT-specific data modalities such as wireless data, drone images, and smart home sensor data (e.g., motion, energy, humidity, temperature) that have not been explored in existing FL benchmarks. Moreover, these datasets target some of the most representative applications and innovative use cases of AIoT that are not possible with other technologies.
To facilitate the community benchmark the performance of the datasets and ensure reproducibility, FedAIoT includes a unified end-to-end FL framework for AIoT, which covers the complete FL-for-AIoT pipeline: from non-independent and identically distributed (non-IID) data partitioning, IoT-specific data preprocessing, to IoT-friendly models, FL hyperparameters, and IoT-factor emulator. Our framework also includes the implementations of popular schemes, models, and techniques involved in each stage of the FL-for-AIoT pipeline.
@@
We have conducted a systematic benchmarking on the eight datasets using the end-to-end framework. Specifically, we examine the impact of varying degrees of non-IID data distributions, FL optimizers, and client sampling ratios on the performance of FL. We also evaluate the impact of noisy labels, a prevalent challenge in IoT datasets, as well as the effects of quantized training, a technique that tackles the practical limitation of resource-constrained IoT devices. Our benchmark results provide valuable insights into both the opportunities and challenges of FL for AIoT. Given the significant relevance of FL in the realm of AIoT, we hope FedAIoT could act as a valuable tool to foster advancements in the important area of FL for AIoT. In summary, our work makes the following contributions.
Table 1: Comparison between FedAIoT with existing FL benchmarks.
@@
  • Unified End-to-End FL Framework for AIoT. We introduce the first unified end-to-end FL framework for AIoT that covers the complete FL-for-AIoT pipeline: from non-IID data partitioning, IoT-specific data preprocessing, to IoT-friendly models, FL hyperparameters, and IoT-factor emulator.
  • Novel Design of Noisy Labels. Our benchmark also introduces a novel way to design noisy labels. Real-world FL deployments on IoT devices often encounter data labeling errors, which act as noises in the federated training. To emulate such noises, different from prior benchmarks (Feng et al., 2023; Zhang et al., 2023a) which adopt uniform error distributions, we have designed a new probabilistic label error scheme based on the insight that labels that are similar to each other are more likely to be mislabeled. This design allows us to simulate data label noises more realistically and to benchmark FL algorithms for their robustness to such noises.
  • Quantized Training. Lastly, we are also the first FL benchmark to show the effect of quantized training on both the server and client sides in the context of FL. Previous benchmark like FLUTE (Dimitriadis et al., 2022) demonstrated the effect of quantized training during server-side aggregation for the purpose of communication reduction. In contrast, our benchmark also incorporates quantized training during client-side training to reduce the memory demands of FL on the device. This is a key difference as IoT devices are not just restrained in communication bandwidth but also in on-device memory.
@@ 2 Related Work
The importance of data to FL research pushes the development of FL benchmarks on a variety of data modalities. Existing FL benchmarks, however, predominantly center around curating FL datasets in the domain of computer vision (CV) (He et al., 2021b; Dimitriadis et al., 2022), natural language processing (NLP) Lin et al. (2021); Dimitriadis et al. (2022), medical imaging (Terrail et al., 2022), speech and audio (Zhang et al., 2023a; Dimitriadis et al., 2022), and graph neural networks (He et al., 2021a). For example, FedCV He et al. (2021b), FedNLP (Lin et al., 2021), and FedAudio (Zhang et al., 2023a) focuses on benchmarking CV, NLP, and audio-related datasets and tasks respectively; FLUTE (Dimitriadis et al., 2022) covers a mix of datasets from CV, NLP, and audio; FLamby (Terrail et al., 2022) primarily focuses on medical imaging datasets; and FedMultimodal (Feng et al., 2023) includes multimodal datasets in the domain of emotion recognition, healthcare, multimedia, and social media. As summarized in Table 1, although these benchmarks have contributed to FL research, a dedicated FL benchmark explicitly tailored for IoT data is absent. Compared to these existing FL benchmarks, FedAIoT is specifically designed to fill this critical gap by providing a dedicated FL benchmark on data collected from a wide range of authentic IoT devices.
@@ 3.1 Datasets
Objective: The objective of FedAIoT is to provide a benchmark that consists of high-quality and well-validated datasets collected from a wide range of IoT devices, sensor modalities, and applications. We have made significant efforts to validate a much larger pool of existing datasets and identified the high-quality ones to include in our benchmark. Table 2 provides an overview of the eight datasets included in FedAIoT. These datasets have diverse sizes (small: less than 7k samples; medium: 11k to 16k samples; and large: more than 60k samples). The rationale behind this design choice is to accommodate researchers who have different computing resources. For example, researchers who have very limited computing resources can still pick the relatively small datasets included in our benchmark to develop and evaluate their algorithms. In this section, we provide a brief overview of each included dataset.
@@
WISDM: The Wireless Sensor Data Mining (WISDM) dataset (Weiss et al., 2019; Lockhart et al., 2011) is one of the widely used datasets for the task of daily activity recognition using accelerometer and gyroscope sensor data collected from smartphones and smartwatches. WISDM includes data collected from $51$ participants performing $18$ daily activities, each in a $3$-minute session. We combined activities such as eating soup, chips, pasta, and sandwiches into a single category called “eating”, and removed uncommon activities related to playing with balls, such as kicking, catching, or dribbling. We randomly selected $45$ participants as the training set and the remaining six participants were assigned to the test set. Given that the smartwatch data and smartphone data were not collected simultaneously for most subjects and thus were not synchronized precisely, we partition WISDM into two independent datasets: WISDM-W with smartwatch data only and WISDM-P with smartphone data only. The total number of samples in the training and test set is $16,569$ and $4,103$ for WISDM-W and $13,714$ and $4,073$ for WISDM-P respectively. No Licence was explicitly mentioned on the dataset homepage.
@@
UT-HAR: The UT-HAR dataset (Yousefi et al., 2017) is a Wi-Fi dataset for the task of contactless activity recognition. The Wi-Fi data are in the form of Channel State Information (CSI) collected using three pairs of antennas and an Intel 5300 Network Interface Card (NIC), with each antenna pair capable of capturing $30$ subcarriers of CSI. UT-HAR comprises data collected from participants performing seven activities such as walking and running. UT-HAR contains a pre-determined training and test set. The total number of training and test samples is $3,977$ and $500$ respectively.
@@
Widar: The Widar dataset (Yang, 2020; Zheng et al., 2019) is designed for contactless gesture recognition using Wi-Fi signal strength measurements collected from strategically placed access points. This system uses an Intel $5300$ NIC with $3\times 3$ antenna pairs. The dataset features data from $17$ participants showcasing $22$ unique gestures, including actions like push, pull, sweeping, and clapping. To maintain consistency between training and test sets, only gestures recorded from more than three users are retained. This decision ensures the test set doesn’t have data from the same participants as the training set. As a result, the balanced dataset includes nine gestures with $11,372$ samples in the training set and $5,222$ in the test set. The dataset is licensed under the Creative Commons Attribution-NonCommercial 4.0 International Licence (CC BY 4).
@@
VisDrone: The VisDrone dataset (Zhu et al., 2021) is a large-scale dataset dedicated to the task of object detection in aerial images captured by drone cameras. VisDrone includes a total of $263$ video clips, which contains $179,264$ frames and $2,512,357$ labeled objects. The labeled objects fall into $12$ categories (e.g., “pedestrian”, “bicycle”, and “car”), recorded under a variety of scenarios such as crowded urban areas, highways, and parks. The dataset contains a pre-determined training and test set. The total number of samples in the training and test set is $6,471$ and $1,610$ respectively. The dataset is licenced under Creative Commons Attribution-NonCommercial-ShareAlike 3.0 License.
@@
CASAS: The CASAS dataset (Schmitter-Edgecombe and Cook, 2009), derived from the CASAS smart home project, is a smart home sensor dataset for the task of recognizing activities of daily living (ADL) based on sequences of sensor states over time to support the application of independent living. Data were collected from three distinct apartments, each equipped with three types of sensors: motion sensors, temperature sensors, and door sensors. We have selected five specific datasets from CASAS named “Milan”, “Cairo”, “Kyoto2”, “Kyoto3”, and “Kyoto4” based on the uniformity of their sensor data representation. The original ADL categories within each dataset have been consolidated into $11$ categories related to home activities such as “sleep”, “eat”, and “bath”. Activities not fitting within these categories were collectively classified as “other”. The training and test set was made using an 80-20 split. Each data sample is a categorical time series of length $2,000$, representing sensor states over a certain period of time. The total number of samples in the training and test set is $12,190$ and $3,048$ respectively. No Licence was mentioned on the dataset homepage.
@@
EPIC-SOUNDS: The EPIC-SOUNDS dataset (Huh et al., 2023) is a large-scale collection of audio recordings for the task of audio-based human activity recognition for augmented reality applications. The audio data were collected from a head-mounted microphone, containing more than $100,000$ categorized segments distributed across $44$ distinct classes. The dataset contains a pre-determined training and test set. The total number of training and test samples is $60,055$ and $40,175$ respectively. The dataset is under CC BY 4 Licence.
@@ 3.2.1 Non-IID Data Partitioning
A key characteristic of FL is that data distribution at different clients is non-IID. The objective of non-IID data partitioning is to partition the training set such that data allocated to different clients follow the non-IID distribution. The eight datasets included in FedAIoT cover three fundamental tasks: classification, regression, and object detection. As summarized in Table 3, FedAIoT incorporates three different non-IID data partitioning schemes that are designed for the three tasks respectively.
@@
Scheme#1: Non-IID Partition over Output Labels. For the task of classification (WISDM-W, WISDM-P, UT-HAR, Widar, CASAS, EPIC-SOUNDS) with $C$ classes, we first generate a distribution over the classes for each client by drawing from a Dirichlet distribution (formal definition in Appendix A.2) with parameter $\alpha$ (Hsu et al., 2019), where lower values of $\alpha$ generate more skewed distribution favoring a few classes whereas higher values of $\alpha$ result in more balanced class distributions. We use the same $\alpha$ to determine the number of samples each client receives. In addition, by drawing from a Dirichlet distribution with parameter $\alpha$, we create a distribution over the total number of samples, which is then used to allocate a varying number of samples to each client, where lower values of $\alpha$ lead to a few clients holding a majority of the samples whereas higher values of $\alpha$ create a more balanced distribution of samples across clients. Therefore, this approach allows us to generate non-IID data partitions such that both class distribution and the number of samples can vary across the clients.
Scheme#2: Non-IID Partition over Input Features. The task of object detection (VisDrone) does not have specific classes. In such case, we use the input features to create non-IID partitions. Specifically, similar to He et al. (2021b), we first used ImageNet (Russakovsky et al., 2015) features generated from the VGG19 model (Liu and Deng, 2015), which encapsulate visual information required for subsequent analysis. With these ImageNet features as inputs, we performed clustering in the feature space using $K$-nearest neighbors to partition the dataset into clusters. Each cluster is a pseudo-class, representing a set of images sharing common visual characteristics as per the extracted ImageNet features. Lastly, Dirichlet allocation was applied on top of the pseudo-classes to create the non-IID distribution across different clients.
Scheme#3: Non-IID Partition over Output Distribution. For the task of regression (AEP) where output is characterized as a continuous variable, we utilize Quantile Binning (Pyle, 1999). Specifically, we divide the range of the output variable into equal groups or quantiles, ensuring that each bin accommodates roughly the same number of samples. Each category or bin is treated as a pseudo-class. Note that the number of quantiles can be set to any value in our framework. We used ten as an example to demonstrate the results. Lastly, we apply Dirichlet allocation to generate the non-IID distribution of data across the clients.

### v3

@@ Abstract
There is a significant relevance of federated learning (FL) in the realm of Artificial Intelligence of Things (AIoT). However, most of existing FL works are not conducted on datasets collected from authentic IoT devices that capture unique modalities and inherent challenges of IoT data. To fill this critical gap, in this work, we introduce FedAIoT, an FL benchmark for AIoT. FedAIoT includes eight datasets collected from a wide range of IoT devices. These datasets cover unique IoT modalities and target representative applications of AIoT. FedAIoT also includes a unified end-to-end FL framework for AIoT that simplifies benchmarking the performance of the datasets. Our benchmark results shed light on the opportunities and challenges of FL for AIoT. We hope FedAIoT could serve as an invaluable resource to foster advancements in the important field of FL for AIoT. The repository of FedAIoT is maintained at https://github.com/AIoT-MLSys-Lab/FedAIoT.
@@ 1 Introduction
Internet of Things (IoT) such as smartphones, drones, and sensors deployed at homes are ubiquitous today. The advances in Artificial Intelligence (AI) have boosted the integration of IoT and AI that turns Artificial Intelligence of Things (AIoT) into reality. However, data captured by IoT devices usually contain privacy-sensitive information. In recent years, federated learning (FL) has emerged as a privacy-preserving solution that allows the extraction of knowledge from collected data while keeping the data solely on the devices (Kairouz et al., 2021; Wang et al., 2021; Zhang et al., 2022).
Despite the significant relevance of FL in the AIoT realm, as summarized in Table 9 in Appendix A, most existing FL works are conducted on well-known datasets such as CIFAR-10 and CIFAR-100. These datasets, however, do not originate from authentic IoT devices and thus fail to capture the unique modalities and inherent challenges associated with real-world IoT data. This discrepancy underscores a strong need for an IoT-oriented FL benchmark to fill this critical gap.
In this work, we present FedAIoT, an FL benchmark for AIoT (Figure 1). At its core, FedAIoT includes eight well-chosen datasets collected from a wide range of IoT devices from smartphones, smartwatches, and Wi-Fi routers, to drones, smart home sensors, and head-mounted devices that either have already become an indispensable part of people’s daily lives or are driving emerging applications. These datasets encapsulate a variety of unique IoT-specific data modalities such as wireless data, drone images, and smart home sensor data (e.g., motion, energy, humidity, temperature) that have not been explored in existing FL benchmarks. Moreover, these datasets target some of the most representative applications and innovative use cases of AIoT.
To facilitate benchmarking and ensure reproducibility, FedAIoT includes a unified end-to-end FL framework for AIoT, which covers the complete FL-for-AIoT pipeline: from non-independent and identically distributed (non-IID) data partitioning, IoT-specific data preprocessing, to IoT-friendly models, FL hyperparameters, and IoT-factor emulator. Our framework also includes the implementations of popular schemes, models, and techniques involved in each stage of the FL-for-AIoT pipeline.
@@
We have conducted systematic benchmarking on the eight datasets using the end-to-end framework. Specifically, we examine the impact of varying degrees of non-IID data distributions, FL optimizers, and client sampling ratios on the performance of FL. We also evaluate the impact of noisy labels, a prevalent challenge in IoT datasets, as well as the effects of quantized training, a technique that tackles the practical limitation of resource-constrained IoT devices. Our benchmark results provide valuable information about both the opportunities and challenges of FL for AIoT. Given the significant relevance of FL in the realm of AIoT, we hope FedAIoT could act as a valuable tool to promote advancements in the important area of FL for AIoT. In summary, our work makes the following contributions.
Table 1: Comparison between FedAIoT and existing FL benchmarks.
@@
  • Unified End-to-End FL Framework for AIoT. We introduce the first unified end-to-end FL framework for AIoT that covers the complete FL-for-AIoT pipeline: from non-IID data partitioning, and IoT-specific data preprocessing, to IoT-friendly models, FL hyperparameters, and IoT-factor emulators.
  • Novel Design of Noisy Labels. Our benchmark also introduces a novel way to design noisy labels. Real-world FL deployments on IoT devices often encounter data labeling errors, which act as noises in federated training. To emulate such noises, different from prior benchmarks that adopt either uniform error distribution or non-uniform error distributions with the assumption that one label can only be mislabeled as another specific label with a random probability, we have designed a new label transition probability matrix based on the insight that labels that are similar to each other are more likely to be mislabeled. This design allows us to simulate data label noises with more realistic transition probabilities and to benchmark FL algorithms for their robustness to such noises.
  • Quantized Training. Lastly, we are also the first FL benchmark to show the effect of quantized training on both server and client sides in the context of FL. Previous benchmarks such as FLUTE (Garcia et al., 2022) demonstrate the effect of quantized training during server-side aggregation for communication reduction. In contrast, our benchmark also incorporates quantized training during client-side training to reduce the memory demands of FL on the device. This is a key difference as IoT devices are often limited by not onlycommunication bandwidth but also on-device memory capacity.
@@ 2 Related Work
The importance of data to FL research pushes the development of FL benchmarks on a variety of data modalities. Existing FL benchmarks, however, predominantly center around curating FL datasets in the domain of computer vision (CV) (He et al., 2021b; Garcia et al., 2022), natural language processing (NLP) Lin et al. (2022); Garcia et al. (2022), medical imaging (du Terrail et al., 2022), speech and audio (Zhang et al., 2023; Garcia et al., 2022), and graph neural networks (He et al., 2021a). For example, FedCV He et al. (2021b), FedNLP (Lin et al., 2022), and FedAudio (Zhang et al., 2023) focuses on benchmarking CV, NLP, and audio-related datasets and tasks respectively; FLUTE (Garcia et al., 2022) covers a mix of datasets from CV, NLP, and audio; FLamby (du Terrail et al., 2022) primarily focuses on medical images; and FedMultimodal (Feng et al., 2023) includes multimodal datasets in the domain of emotion recognition, healthcare, multimedia, and social media. As summarized in Table 1, although these benchmarks have contributed to FL research, a dedicated FL benchmark explicitly tailored for IoT data is absent. Compared to these existing FL benchmarks, FedAIoT is specifically designed to fill this critical gap by providing a dedicated FL benchmark on data collected from a wide range of authentic IoT devices.
@@ 3.1 Datasets
Objective: The objective of FedAIoT is to provide a benchmark consisting of well-validated and high-quality datasets collected from a wide range of IoT devices, sensor modalities, and applications. We have made significant efforts to examine a much larger pool of existing datasets and select the high-quality ones to include in our benchmark. Table 2 provides an overview of the eight high-quality datasets included in FedAIoT. These datasets have diverse sizes (small: less than 7k samples; medium: 11k to 16k samples; and large: more than 60k samples). The rationale behind this design choice is to accommodate researchers with different computing resources. For example, researchers with limited computing resources can still pick the relatively small datasets in our benchmark to develop and evaluate their algorithms. In this section, we provide a brief overview of each included dataset. More details about the datasets are provided in Appendix B and Table 10.
@@
WISDM: The Wireless Sensor Data Mining (WISDM) dataset (Weiss et al., 2019; Lockhart et al., 2011) is one of the widely used datasets for the task of daily activity recognition using accelerometer and gyroscope sensor data collected from smartphones and smartwatches. WISDM includes data collected from $51$ subjects performing $18$ daily activities, each in a $3$-minute session. We combined activities such as eating soup, chips, pasta, and sandwiches into a single category called “eating”, and removed uncommon activities related to playing with balls, such as kicking, catching, or dribbling. We randomly selected $45$ subjects as the training set and the remaining six subjects were assigned to the test set. Given that the smartwatch data and smartphone data were not collected simultaneously for most subjects and thus were not synchronized precisely, we partition WISDM into two independent datasets: WISDM-W with smartwatch data only and WISDM-P with smartphone data only. The total number of samples in the training and test set is $16,569$ and $4,103$ for WISDM-W and $13,714$ and $4,073$ for WISDM-P respectively. No Licence was explicitly mentioned on the dataset homepage.
@@
UT-HAR: The UT-HAR dataset (Yousefi et al., 2017) is a dataset for contactless activity recognition based on Wi-Fi signals. The Wi-Fi data are in the form of Channel State Information (CSI) collected using three pairs of antennas and an Intel 5300 Network Interface Card (NIC), with each antenna pair capable of capturing $30$ subcarriers of CSI. UT-HAR comprises data collected from subjects performing seven activities such as walking and running. UT-HAR contains a pre-determined training and test set. The total number of training and test samples is $3,977$ and $500$ respectively. No Licence was mentioned on the dataset homepage.
@@
Widar: The Widar dataset (Yang, 2020; Zheng et al., 2019) is designed for contactless gesture recognition using Wi-Fi signal strength measurements collected from strategically placed access points. The data collection system uses an Intel $5300$ NIC with $3\times 3$ antenna pairs. The dataset includes data from $17$ subjects performing $22$ unique gestures like push, pull, sweeping, and clapping. There are some gesture classes that were only contributed by one subject. Including those would risk the model overfitting for that subject instead of the gesture itself. As such, only gestures recorded by more than three subjects are retained, with data from two subjects used during training and the remaining one for the test set. This decision ensures that the test set does not have data from the same subjects as the training set. The resulting dataset includes nine gestures with $11,372$ samples in the training set and $5,222$ in the test set. The dataset is licensed under the Creative Commons Attribution-NonCommercial 4.0 International Licence (CC BY 4).
@@
VisDrone: The VisDrone dataset (Zhu et al., 2021) is a large-scale dataset dedicated to object detection in aerial images captured by drone cameras. VisDrone includes a total of $263$ video clips, which contain $179,264$ frames and $2,512,357$ labeled objects. The labeled objects fall into $12$ categories (e.g., “pedestrian”, “bicycle”, and “car”), recorded under various scenarios such as crowded urban areas, highways, and parks. The dataset contains a pre-determined training and test set. The total number of samples in the training and test set is $6,471$ and $1,610$ respectively. The dataset is licensed under Creative Commons Attribution-NonCommercial-ShareAlike 3.0 License.
@@
CASAS: The CASAS dataset (Schmitter-Edgecombe and Cook, 2009), derived from the CASAS smart home project, is a smart home sensor dataset for the task of recognizing activities of daily living (ADL) based on sequences of sensor states over time to support the application of independent living. Data were collected from three distinct apartments, each equipped with three types of sensors: motion sensors, temperature sensors, and door sensors. We have selected five specific datasets from CASAS named “Milan”, “Cairo”, “Kyoto2”, “Kyoto3”, and “Kyoto4” based on the uniformity of their sensor data representation. The original ADL categories within each dataset have been consolidated into $11$ categories related to home activities such as “sleep”, “eat”, and “bath”. Activities not fitting within these categories were collectively classified as “other”. The training and test set was made using an 80-20 split. Each data sample is a categorical time series of length $2,000$, representing sensor states over a certain period. The total number of samples in the training and test set is $12,190$ and $3,048$ respectively. No Licence was mentioned on the dataset homepage.
@@
EPIC-SOUNDS: The EPIC-SOUNDS dataset (Huh et al., 2023) is a large-scale collection of audio recordings for audio-based human activity recognition for augmented reality applications. The audio data were collected from a head-mounted microphone, containing more than $100,000$ categorized segments distributed across $44$ distinct classes. The dataset contains a pre-determined training and test set. The total number of training and test samples is $60,055$ and $40,175$ respectively. The dataset is under CC BY 4 Licence.
@@ 3.2.1 Non-IID Data Partitioning
A key characteristic of FL is that data distribution at different clients is non-IID. The objective of non-IID data partitioning is to partition the training set such that data allocated to different clients follow the non-IID distribution. The eight datasets included in FedAIoT cover three fundamental tasks: classification, regression, and object detection. As summarized in Table 3, FedAIoT incorporates three different non-IID data partitioning schemes designed for the three tasks respectively.
@@
Scheme#1: Non-IID Partition over Output Labels. For the task of classification (WISDM-W, WISDM-P, UT-HAR, Widar, CASAS, EPIC-SOUNDS) with $C$ classes, we first generate a distribution over the classes for each client by drawing from a Dirichlet distribution (formal definition in Appendix C) with parameter $\alpha$ (Hsu et al., 2019), where lower values of $\alpha$ generate more skewed distribution whereas higher values of $\alpha$ result in more balanced class distributions. We use the same $\alpha$ to determine the number of samples each client contains. In addition, by drawing from a Dirichlet distribution with parameter $\alpha$, we create a distribution over the total number of samples, which is then used to allocate a varying number of samples to each client, where lower values of $\alpha$ lead to a few clients holding a majority of the samples whereas higher values of $\alpha$ create a more balanced distribution of samples across clients. Therefore, this approach allows us to generate non-IID data partitions such that both class distribution and the number of samples can vary across the clients.
Scheme#2: Non-IID Partition over Input Features. The task of object detection (VisDrone) does not have specific classes. In such case, we use the input features to create non-IID partitions. Specifically, similar to He et al. (2021b), we first used ImageNet (Russakovsky et al., 2015) features generated from the VGG19 model (Liu and Deng, 2015), which encapsulate visual information required for subsequent analysis. With these ImageNet features as inputs, we performed clustering in the feature space using $K$-nearest neighbors to partition the dataset into clusters. Each cluster is a pseudo-class, representing images sharing common visual characteristics as per the extracted ImageNet features. Lastly, Dirichlet allocation was applied on top of the pseudo-classes to create the non-IID distribution across different clients.
Scheme#3: Non-IID Partition over Output Distribution. For the regression task (AEP) where output is characterized as a continuous variable, we utilize Quantile Binning (Pyle, 1999). Specifically, we divide the range of the output variable into equal groups or quantiles, ensuring that each bin accommodates roughly the same number of samples. Each category or bin is treated as a pseudo-class. Note that the number of quantiles can be set to any value in our framework. We used ten as an example to demonstrate the results. Lastly, we apply Dirichlet allocation to generate the non-IID distribution of data across the clients.